OpenStudy (anonymous):

Solve the inequality: -3-4(2x-5)+10=-3(2x+6)-10

OpenStudy (anonymous):

-3-8x+20+10=-6x-18-10 27-8x=-6x-28 -2x=-55 x=-55/2